A ball is drawn from a box containing 6 red balls, 4 white balls and 5 black balls . What is the probability that it is not red ball?



1
Expert's answer
2022-05-15T23:08:16-0400

The total number of outcomes:

"N=6+4+5=15."

The number of favorable outcomes (the drawn ball is white or black):

"n=4+5=9."

The probability that the drawn ball is not red ball:

"P=\cfrac{n}{N}=\cfrac{9}{15}=\cfrac{3}{5}=0.6."